# Calibrating Deep Neural Network using Euclidean Distance

### v1

@@ A.4. Convexity of the Sum of Strictly Convex and Convex Functions
(42) $f\big{(}\theta x+(1-\theta)y\big{)}<\theta f(x)+(1-\theta)f(y).$
@@
(43) $g\big{(}\theta x+(1-\theta)y\big{)}\leq\theta g(x)+(1-\theta)g(y).$
@@
(44) $\begin{split}h\big{(}\theta x+(1-\theta)y\big{)}&=f\big{(}\theta x+(1-\theta)y\big{)}+g\big{(}\theta x+(1-\theta)y\big{)}\\ &<\theta f(x)+(1-\theta)f(y)+\theta g(x)+(1-\theta)g(y)\\ &=\theta\big{(}f(x)+g(x)\big{)}+(1-\theta)\big{(}f(y)+g(y)\big{)}\\ &=\theta h(x)+(1-\theta)h(y).\end{split}$
@@
A.5. Proof that the Focal Calibration Loss Satisfies the Strictly Order-Preserving Property
Theorem 2.
@@ A.5. Proof that the Focal Calibration Loss Satisfies the Strictly Order-Preserving Property
From Def. 4, we aim to show that the Focal Calibration Loss (FCL) satisfies:
(45) $q_{i}^{*}<q_{j}^{*}\Rightarrow\eta_{i}<\eta_{j},$
where $\bm{q}^{*}=[q_{1}^{*},q_{2}^{*},\dots,q_{K}^{*}]^{\top}$ is the minimizer of the pointwise conditional risk $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\bm{q})$.
@@
The Focal Calibration Loss for a single sample $\bm{x}$ with true label $y\in\{1,2,\ldots,K\}$ is defined as:
(46) $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q},y)=-(1-q_{y})^{\gamma}\log q_{y}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2},$
@@
  • $\bm{q}=[q_{1},q_{2},\dots,q_{K}]^{\top}$ is the predicted probability distribution over the $K$ classes.
  • $\eta_{i}=\mathbb{P}(y=i\mid\bm{x})$ is the true class-posterior probability for class $i$.
@@
The pointwise conditional risk $\mathcal{R}^{\mathcal{L}}(\bm{q})$ for the loss function $\mathcal{L}$ is defined as:
(47) $\mathcal{R}^{\mathcal{L}}(\bm{q})=\mathbb{E}_{y\mid\bm{x}}[\mathcal{L}(\bm{q},y)]=\sum_{i=1}^{K}\eta_{i}\mathcal{L}(\bm{q},i).$
@@
(48) $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\bm{q})=\sum_{i=1}^{K}\eta_{i}\left[-(1-q_{i})^{\gamma}\log q_{i}\right]+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}.$
@@
Our objective is to find the minimizer $\bm{q}^{*}$ of $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\bm{q})$:
(49) $\bm{q}^{*}=\arg\min_{\bm{q}}\left\{\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\bm{q})\mid\sum_{i=1}^{K}q_{i}=1,\quad q_{i}>0\quad\forall i\right\}.$
@@
(50) $\mathcal{L}(\bm{q},\mu)=\sum_{i=1}^{K}\eta_{i}\left[-(1-q_{i})^{\gamma}\log q_{i}\right]+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}+\mu\left(\sum_{i=1}^{K}q_{i}-1\right).$
@@
For each $i$, compute the partial derivative of $\mathcal{L}(\bm{q},\mu)$ with respect to $q_{i}$:
(51) $\frac{\partial\mathcal{L}}{\partial q_{i}}=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i})+\mu.$
Set the derivative to zero at the minimizer $\bm{q}^{*}$:
(52) $\frac{\partial\mathcal{L}}{\partial q_{i}}\bigg{|}_{\bm{q}=\bm{q}^{*}}=0.$
@@
(53) $\eta_{i}\left[\gamma(1-q_{i}^{*})^{\gamma-1}\log q_{i}^{*}+\frac{(1-q_{i}^{*})^{\gamma}}{q_{i}^{*}}\right]+2\lambda(q_{i}^{*}-\eta_{i})+\mu=0.$
@@
(55) $\Phi(q_{i}^{*})=\gamma(1-q_{i}^{*})^{\gamma-1}\log q_{i}^{*}+\frac{(1-q_{i}^{*})^{\gamma}}{q_{i}^{*}}.$
@@
(56) $\eta_{i}\Phi(q_{i}^{*})+2\lambda(q_{i}^{*}-\eta_{i})=-\mu\quad\forall i.$
@@
(58) $\eta_{i}\left[\Phi(q_{i}^{*})-\Phi(q_{j}^{*})\right]+2\lambda(q_{i}^{*}-q_{j}^{*})=0.$
@@
(59) $\left[\eta_{i}\Phi(q_{i}^{*})-\eta_{j}\Phi(q_{j}^{*})\right]+2\lambda\left(q_{i}^{*}-\eta_{i}-q_{j}^{*}+\eta_{j}\right)=0.$
@@
A.6. Properties of $\sigma^{\gamma,\lambda}$
Lemma 3.
@@ A.6. Properties of $\sigma^{\gamma,\lambda}$
(60) $\sigma^{\gamma,\lambda}(q)=(1-q)^{\gamma}-\gamma q\log q\,(1-q)^{\gamma-1}-2\lambda q,$
@@
  • $-\gamma q\log q\,(1-q)^{\gamma-1}$ is continuous and differentiable since $\log q$ is continuous on $q\in(0,1)$.
  • $-2\lambda q$ is a linear function and thus continuous and differentiable.
@@
As $q\to 0^{+}$, we have $(1-q)^{\gamma}\to 1$, $q\log q\to 0$, and $-2\lambda q\to 0$.
(61) $\sigma^{\gamma,\lambda}(q)\to 1-0+0=1.$
@@
As $q\to 1^{-}$, $(1-q)^{\gamma}\to 0$, $-\gamma q\log q\,(1-q)^{\gamma-1}\to 0$, and $-2\lambda q\to-2\lambda$. Therefore:
(62) $\sigma^{\gamma,\lambda}(q)\to 0+0-2\lambda=-2\lambda.$
@@
(63) $\begin{split}\sigma^{\prime\gamma,\lambda}(q)&=\frac{d}{dq}\left[(1-q)^{\gamma}-\gamma q\log q\,(1-q)^{\gamma-1}-2\lambda q\right]\\ &=-\gamma(1-q)^{\gamma-1}-\gamma\log q\,(1-q)^{\gamma-1}+\gamma q\frac{1}{q}(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=-\gamma(1-q)^{\gamma-1}-\gamma\log q\,(1-q)^{\gamma-1}+\gamma(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=\gamma\left[-(1-q)^{\gamma-1}+(1-q)^{\gamma-1}\right]-\gamma\log q\,(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=-\gamma\log q\,(1-q)^{\gamma-1}+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda.\end{split}$
@@
  • If $10\leq\gamma<$ 1, then $(\gamma-1)<0$, but $\log q<0$, so the product is positive; however, $q(1-q)^{\gamma-2}$ remains finite and positive.
@@
Figure 6. Plot of $\sigma^{\gamma,\lambda}(q)$ vs. $q$ for different $\gamma$ and $\lambda$
[…]
This Plot demonstrate that $\sigma^{2,1}(q)$ decreases from positive to negative as $q$ increases, confirming the theoretical properties established in the lemma.
@@
A.7. Proof of Thm. (6): The calibration loss defined as the mean squared $L_{2}$ norm is a strictly proper scoring rule for both binary and multiclass classification.
@@
In the proof, $Y\sim P$ denotes a random variable $Y$ that follows the probability distribution $P$. The predicted probability for class $k$ is $q_{k}$, and the true probability for class $k$ is $p_{k}$. The expectation $\mathbb{E}_{Y\sim P}[S(Q,Y)]$ represents the expected value of the scoring function $S(Q,Y)$ over the distribution $P$. The calibration loss $\mathcal{L}_{\text{calib}}$ is defined as the mean squared $L_{2}$ norm, and it measures the squared differences between predicted probabilities $\hat{p}_{i}$ and actual binary outcomes $y_{i}$ for binary classification, or between predicted probability vectors $\hat{\bm{p}}_{i}$ and true outcome vectors $\bm{y}_{i}$ for multiclass classification. The derivatives $\frac{\partial}{\partial q_{j}}$ and $\frac{\partial^{2}}{\partial q_{j}^{2}}$ denote the first and second partial derivatives w.r.t. the predicted probability $q_{j}$, used to find and verify the maxima of the expected score.
@@
(68) $\begin{split}\frac{\partial}{\partial q}\mathbb{E}_{Y\sim P}[S(Q,Y)]&=-2p(q-1)-2(1-p)q\\ &=-2pq+2p-2q+2pq\\ &=2p-2q\Rightarrow 0=2p-2q\Rightarrow p=q\end{split}$
@@
(69) $\frac{\partial^{2}}{\partial q^{2}}\mathbb{E}_{Y\sim P}[S(Q,Y)]=-2<0$
@@
(70) $\mathbb{E}_{Y\sim P}[S(Q,Y)]$ $=\sum_{k}P(Y=k)\left[-\sum_{i}(q_{i}-y_{i})^{2}\right]$
@@
(72) $\frac{\partial}{\partial q_{j}}\mathbb{E}_{Y\sim P}[S(Q,Y)]$ $=2p_{j}-2q_{j}-2p_{K}+2\left(1-\sum_{i\neq K}q_{i}\right)$
@@
(74) $\frac{\partial^{2}}{\partial q_{j}^{2}}\mathbb{E}_{Y\sim P}[S(Q,Y)]=-2-2=-4<0$
@@
A.8. Proof that the Focal Calibration Loss is Strictly Proper

### v2

@@ A.4 Convexity of the Sum of Strictly Convex and Convex Functions
$f\big{(}\theta x+(1-\theta)y\big{)}<\theta f(x)+(1-\theta)f(y).$ (41)
@@
$g\big{(}\theta x+(1-\theta)y\big{)}\leq\theta g(x)+(1-\theta)g(y).$ (42)
@@
$\begin{split}h\big{(}\theta x+(1-\theta)y\big{)}&=f\big{(}\theta x+(1-\theta)y\big{)}+g\big{(}\theta x+(1-\theta)y\big{)}\\ &<\theta f(x)+(1-\theta)f(y)+\theta g(x)+(1-\theta)g(y)\\ &=\theta\big{(}f(x)+g(x)\big{)}+(1-\theta)\big{(}f(y)+g(y)\big{)}\\ &=\theta h(x)+(1-\theta)h(y).\end{split}$ (43)
@@
A.5 Proof that the Focal Calibration Loss Satisfies the Strictly Order-Preserving Property
Theorem A.2.
@@ A.5 Proof that the Focal Calibration Loss Satisfies the Strictly Order-Preserving Property
From Def. LABEL:def:Strictly_Order-Preserving_Property, we aim to show that the Focal Calibration Loss (FCL) satisfies:
$q_{i}^{*}<q_{j}^{*}\Rightarrow\eta_{i}<\eta_{j},$ (44)
where $\boldsymbol{q}^{*}=[q_{1}^{*},q_{2}^{*},\dots,q_{K}^{*}]^{\top}$ is the minimizer of the pointwise conditional risk $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})$.
@@
The Focal Calibration Loss for a single sample $\boldsymbol{x}$ with true label $y\in\{1,2,\ldots,K\}$ is defined as:
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q},y)=-(1-q_{y})^{\gamma}\log q_{y}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2},$ (45)
@@
  • $\boldsymbol{q}=[q_{1},q_{2},\dots,q_{K}]^{\top}$ is the predicted probability distribution over the $K$ classes.
  • $\eta_{i}=\mathbb{P}(y=i\mid\boldsymbol{x})$ is the true class-posterior probability for class $i$.
@@
The pointwise conditional risk $\mathcal{R}^{\mathcal{L}}(\boldsymbol{q})$ for the loss function $\mathcal{L}$ is defined as:
$\mathcal{R}^{\mathcal{L}}(\boldsymbol{q})=\mathbb{E}_{y\mid\boldsymbol{x}}[\mathcal{L}(\boldsymbol{q},y)]=\sum_{i=1}^{K}\eta_{i}\mathcal{L}(\boldsymbol{q},i).$ (46)
@@
$\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})=\sum_{i=1}^{K}\eta_{i}\left[-(1-q_{i})^{\gamma}\log q_{i}\right]+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}.$ (47)
@@
Our objective is to find the minimizer $\boldsymbol{q}^{*}$ of $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})$:
$\boldsymbol{q}^{*}=\arg\min_{\boldsymbol{q}}\left\{\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})\mid\sum_{i=1}^{K}q_{i}=1,\quad q_{i}>0\quad\forall i\right\}.$ (48)
@@
$\mathcal{L}(\boldsymbol{q},\mu)=\sum_{i=1}^{K}\eta_{i}\left[-(1-q_{i})^{\gamma}\log q_{i}\right]+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}+\mu\left(\sum_{i=1}^{K}q_{i}-1\right).$ (49)
@@
For each $i$, compute the partial derivative of $\mathcal{L}(\boldsymbol{q},\mu)$ with respect to $q_{i}$:
$\frac{\partial\mathcal{L}}{\partial q_{i}}=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i})+\mu.$ (50)
Set the derivative to zero at the minimizer $\boldsymbol{q}^{*}$:
$\frac{\partial\mathcal{L}}{\partial q_{i}}\bigg{|}_{\boldsymbol{q}=\boldsymbol{q}^{*}}=0.$ (51)
@@
$\eta_{i}\left[\gamma(1-q_{i}^{*})^{\gamma-1}\log q_{i}^{*}+\frac{(1-q_{i}^{*})^{\gamma}}{q_{i}^{*}}\right]+2\lambda(q_{i}^{*}-\eta_{i})+\mu=0.$ (52)
@@
$\Phi(q_{i}^{*})=\gamma(1-q_{i}^{*})^{\gamma-1}\log q_{i}^{*}+\frac{(1-q_{i}^{*})^{\gamma}}{q_{i}^{*}}.$ (54)
@@
$\eta_{i}\Phi(q_{i}^{*})+2\lambda(q_{i}^{*}-\eta_{i})=-\mu\quad\forall i.$ (55)
@@
$\eta_{i}\left[\Phi(q_{i}^{*})-\Phi(q_{j}^{*})\right]+2\lambda(q_{i}^{*}-q_{j}^{*})=0.$ (57)
@@
$\left[\eta_{i}\Phi(q_{i}^{*})-\eta_{j}\Phi(q_{j}^{*})\right]+2\lambda\left(q_{i}^{*}-\eta_{i}-q_{j}^{*}+\eta_{j}\right)=0.$ (58)
@@
A.6 Properties of $\sigma^{\gamma,\lambda}$
Lemma A.3.
@@ A.6 Properties of $\sigma^{\gamma,\lambda}$
$\sigma^{\gamma,\lambda}(q)=(1-q)^{\gamma}-\gamma q\log q\,(1-q)^{\gamma-1}-2\lambda q,$ (59)
@@
  • $\gamma q\log q\,(1-q)^{\gamma-1}$ is continuous and differentiable since $\log q$ is continuous on $q\in(0,1)$.
  • $2\lambda q$ is a linear function and thus continuous and differentiable.
@@
As $q\to 0^{+}$, we have $(1-q)^{\gamma}\to 1$, $q\log q\to 0$, and $2\lambda q\to 0$.
$\sigma^{\gamma,\lambda}(q)\to 1-0+0=1.$ (60)
@@
As $q\to 1^{-}$, $(1-q)^{\gamma}\to 0$, $\gamma q\log q\,(1-q)^{\gamma-1}\to 0$, and $2\lambda q\to-2\lambda$. Therefore:
$\sigma^{\gamma,\lambda}(q)\to 0+0-2\lambda=-2\lambda.$ (61)
@@
$\begin{split}\sigma^{\prime\gamma,\lambda}(q)&=\frac{d}{dq}\left[(1-q)^{\gamma}-\gamma q\log q\,(1-q)^{\gamma-1}-2\lambda q\right]\\ &=-\gamma(1-q)^{\gamma-1}-\gamma\log q\,(1-q)^{\gamma-1}+\gamma q\frac{1}{q}(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=-\gamma(1-q)^{\gamma-1}-\gamma\log q\,(1-q)^{\gamma-1}+\gamma(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=\gamma\left[-(1-q)^{\gamma-1}+(1-q)^{\gamma-1}\right]-\gamma\log q\,(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=-\gamma\log q\,(1-q)^{\gamma-1}+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda.\end{split}$ (62)
@@
  • If $0\leq\gamma<1$, then $(\gamma-1)<0$, but $\log q<0$, so the product is positive; however, $q(1-q)^{\gamma-2}$ remains finite and positive.
@@
Figure 5: Plot of $\sigma^{\gamma,\lambda}(q)$ vs. $q$ for different $\gamma$ and $\lambda$
@@
A.7 Proof: The calibration loss defined as the mean squared $L_{2}$ norm is a strictly proper scoring rule for both binary and multiclass classification.
@@
In the proof, $Y\sim P$ denotes a random variable $Y$ that follows the probability distribution $P$. The predicted probability for class $k$ is $q_{k}$, and the true probability for class $k$ is $p_{k}$. The expectation $\mathbb{E}_{Y\sim P}[S(Q,Y)]$ represents the expected value of the scoring function $S(Q,Y)$ over the distribution $P$. The calibration loss $\mathcal{L}_{\text{calib}}$ is defined as the mean squared $L_{2}$ norm, and it measures the squared differences between predicted probabilities $\hat{p}_{i}$ and actual binary outcomes $y_{i}$ for binary classification, or between predicted probability vectors $\hat{\boldsymbol{p}}_{i}$ and true outcome vectors $\boldsymbol{y}_{i}$ for multiclass classification. The derivatives $\frac{\partial}{\partial q_{j}}$ and $\frac{\partial^{2}}{\partial q_{j}^{2}}$ denote the first and second partial derivatives w.r.t. the predicted probability $q_{j}$, used to find and verify the maxima of the expected score.
@@
$\begin{split}\frac{\partial}{\partial q}\mathbb{E}_{Y\sim P}[S(Q,Y)]&=-2p(q-1)-2(1-p)q\\ &=-2pq+2p-2q+2pq\\ &=2p-2q\Rightarrow 0=2p-2q\Rightarrow p=q\end{split}$ (67)
@@
$\frac{\partial^{2}}{\partial q^{2}}\mathbb{E}_{Y\sim P}[S(Q,Y)]=-2<0$ (68)
@@
$\mathbb{E}_{Y\sim P}[S(Q,Y)]$ $=\sum_{k}P(Y=k)\left[-\sum_{i}(q_{i}-y_{i})^{2}\right]$ (69)
@@
$\frac{\partial}{\partial q_{j}}\mathbb{E}_{Y\sim P}[S(Q,Y)]$ $=2p_{j}-2q_{j}-2p_{K}+2\left(1-\sum_{i\neq K}q_{i}\right)$ (71)
@@
$\frac{\partial^{2}}{\partial q_{j}^{2}}\mathbb{E}_{Y\sim P}[S(Q,Y)]=-2-2=-4<0$ (73)
@@
A.8 Proof that the Focal Calibration Loss is Strictly Proper